# Topologically protected synchronization in networks

### v1

@@ Abstract
and the eigenvalues of $B$ are $\lambda_{1}=k^{(\mathrm{OUT})}-3+4\gamma$ and $\lambda_{2}=k^{(\mathrm{OUT})}-3+2\gamma$. We conclude that the group of $N^{\prime}=3$ TE FC nodes synchronize if $k^{(\mathrm{OUT})}-3+4\gamma<0$, which, on observing that $10<4\gamma<$ 1, amounts to $k^{(\mathrm{OUT})}\leq 2$, or $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$, where we have made use of the fact that, here, $k^{(\mathrm{IN})}=2$. Note that also for the previously seen case $N^{\prime}=2$, where it was $k^{(\mathrm{IN})}=1$, the TC for the synchronization can be written as $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$. In general, for a group of $N^{\prime}$ TE nodes FC, the diagonal elements of the matrix $\bm{B}$ are all equal to $k^{(\mathrm{OUT})}-N^{\prime}(1-\gamma)$, while the off-diagonal ones are in the form $p\gamma$ where $p\in\{1,\ldots,N^{\prime}-2\}$. We have verified that for $N^{\prime}=4$ the TC remains $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$ but for $N^{\prime}=5$ the TC becomes $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}-1$. For details see SM where we also analyze cases where the group forms regular polygons. We stress that these are sufficient conditions; they might be not necessary.
@@
Figure 2: Numerical solution of model (1) with $J=1$ and $N=22$ oscillators on the top of the graph of Fig. 1. Plots correspond to the variables $\theta_{i+1}-\theta_{i}$, for $i=1,\ldots,N-1$. In each group of TE nodes, the initial conditions, as well as the natural frequencies, are chosen very close to each other as to guarantee the strict sufficient conditions explained in the text. It is evident that for several indices $i$, $\theta_{i+1}-\theta_{i}$ drifts away, while there exists a group of indices where $\theta_{i+1}-\theta_{i}$ remains bounded. As can be better checked by the Inset, the bounded variables includes all the five groups of TE nodes and a few others which, however, are manifestly less synchronized than the TE nodes. Note in particular that, unlike the latter, in each group, as analytically predicted, the sign of $\theta_{i+1}-\theta_{i}$ does not change over time.
@@
We illustrate our analysis by solving numerically Numerical_Recipes model (1) in a system of moderate size but involving most of the ideas so far discussed. The graph associated to this system is depicted in Fig. 1. It contains five groups of TE nodes of various kind for which the TC is satisfied. We performed several simulations with different choices of the parameters $\omega_{1},\ldots,\omega_{N}$ and $\theta_{1}(0),\ldots,\theta_{N}(0)$ ranging from situations in which the sufficient conditions for synchronization discussed previously are strictly satisfied, to situations in which only the TC is satisfied. Remarkably, even in these latter situations, the synchronization scenario predicted analytically (via stricter conditions) holds. Figs. 2 and 3 show the behavior of $\theta_{i+1}(t)-\theta_{i}(t)$ in the two situations. Inset of Fig. 3 shows also the rotating numbers $\theta_{i}(t)/t$. For more details see SM .
@@
In conclusion, simulations confirm the local synchronization scenario predicted analytically and extend its region of validity for a quite wider range of initial conditions and spread of natural frequencies: groups of TE nodes with $k^{(\mathrm{OUT})}$ sufficiently smaller than $k^{(\mathrm{IN})}$ (depending on the specific group), synchronize and remain protected from the remainder of the system, which could be even noisy. Moreover, the plots of the rotating numbers strengthen the idea that these groups of TE nodes play the role of independent pacemakers of the system. In fact, in these plots, the number of lines of convergence seems to scale as the number of TE groups. As the coupling $J$ increases, these lines approach each other and eventually converge to a unique line establishing global synchronization, consistently with ArenasPRL ; Note2 . However, for any finite value of $J$, each group keeps its own synchronization status, regardless of the others; this fact seems to hinder global synchronization while allowing for the local one. Of course, this hypothesis needs to be tested in larger systems. Many other questions arise, particularly the following: i) Do real-world networks contain groups of TE nodes? ii) Do real-world networks exist where the number of different groups, say $Q$, scales with the system size $N$? iii) Is it possible to probe efficiently very large systems to quantify $Q$ as well as to identify the nodes of these groups? We anticipate that these three questions meet a positive answer and will be the subject of a subsequent publication.

### v2

@@ Abstract
and the eigenvalues of $B$ are $\lambda_{1}=k^{(\mathrm{OUT})}-3+4\gamma$ and $\lambda_{2}=k^{(\mathrm{OUT})}-3+2\gamma$. We conclude that the group of $N^{\prime}=3$ TE FC nodes synchronize if $k^{(\mathrm{OUT})}-3+4\gamma<0$, which, on observing that $0<4\gamma<1$, amounts to $k^{(\mathrm{OUT})}\leq 2$, or $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$, where we have made use of the fact that, here, $k^{(\mathrm{IN})}=2$. Note that also for the previously seen case $N^{\prime}=2$, where it was $k^{(\mathrm{IN})}=1$, the TC for the synchronization can be written as $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$. In general, for a group of $N^{\prime}$ FC TE nodes, the diagonal elements of the matrix $\bm{B}$ are all equal to $k^{(\mathrm{OUT})}-N^{\prime}(1-\gamma)$, while the off-diagonal ones are in the form $p\gamma$ where $p\in\{1,\ldots,N^{\prime}-2\}$. We have verified that for $N^{\prime}=4$ the TC remains $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$ but for $N^{\prime}=5$ the TC becomes $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}-1$. For details see SM where we also analyze cases where the group forms regular polygons. We stress that these are sufficient conditions; they might be not necessary.
@@
Figure 2: Numerical solution of model (1) with $J=1$ and $N=22$ oscillators on the top of the graph of Fig. 1. Plots correspond to the variables $\theta_{i+1}-\theta_{i}$, for $i=1,\ldots,N-1$. In each group of TE nodes, the initial conditions, as well as the natural frequencies, are chosen very close to each other as to guarantee the strict sufficient conditions explained in the text. It is evident that for several indices $i$, $\theta_{i+1}-\theta_{i}$ drifts away, while there exists a group of indices where $\theta_{i+1}-\theta_{i}$ remains bounded. As can be better checked by the Insets, the bounded variables includes all the five groups of TE nodes and a few others which, however, are manifestly less synchronized than the TE nodes. Note in particular that, unlike the latter, in each group, as analytically predicted, the sign of $\theta_{i+1}-\theta_{i}$ does not change over time. See Fig. S2 and SM for further visualization.
@@
We illustrate our analysis by solving numerically Numerical_Recipes model (1) in a system of moderate size but involving most of the ideas so far discussed. The graph associated to this system is depicted in Fig. 1. It contains five groups of TE nodes of various kind for which the TC is satisfied. We performed several simulations with different choices of the parameters $\omega_{1},\ldots,\omega_{N}$ and $\theta_{1}(0),\ldots,\theta_{N}(0)$ ranging from situations in which the sufficient conditions for synchronization discussed previously are strictly satisfied, to situations in which only the TC is satisfied. Remarkably, even in these latter situations, the synchronization scenario predicted analytically (via stricter conditions) holds. Figs. 2 and 3 show the behavior of $\theta_{i+1}(t)-\theta_{i}(t)$ in the two situations. Inset of Fig. 3 shows also the rotating numbers $\theta_{i}(t)/t$ StrogatzC . For more details see also Fig. S2 and SM .
@@
Figure 4: Particular of Fig. 2. It shows in full detail all the ten phase differences $\theta_{i+1}-\theta_{i}$ associated to the ten links writable in the form $(i,i+1)$ inside the groups of TE nodes of the graph of Fig. 1. It demonstrates the consistency of the general behavior $\theta_{i+1}-\theta_{i}$ with the bounds established for example by Eq. (15), as well as by Eqs. (S11)-(S12) of SM: After an initial exponential drop, $\theta_{i+1}-\theta_{i}$ in each group reaches a smooth regime with a bounded value that is proportional to the spread of the group’s frequencies and inversely proportional to the coupling $J$.
In conclusion, we analytically predict and confirm by simulations the local synchronization of oscillators with heterogeneous frequencies: groups of TE nodes with $k^{(\mathrm{OUT})}$ sufficiently smaller than $k^{(\mathrm{IN})}$ (depending on the group), synchronize and remain protected from the rest of the system, even under noise. Plots of rotating numbers support the idea that TE groups act as independent pacemakers. In fact, the number of convergence lines scales with the number of TE groups and, as the coupling $J$ increases, these lines approach and eventually merge into one, indicating global synchronization, consistently with ArenasPRL ; Pecora . However, for any finite value of $J$, each group of TE oscillators satisfying the TC maintains its own synchronization status, regardless of the others. This is precisely what is expected, e.g., in many biological systems, whose functionality relies on the coexistence of several independent pacemakers. Our exact result could have broad applications since, as shown in MacArthur , in real-world networks, the number of groups of FC TE nodes scales with the system size $N$, ranging from 20 % to 98%.
This work was supported by the National Council for Scientific and Technological Development (CNPq), Brazil, under the Universal Grant (process no. 409180/2023-8).
Supplemental Material for
Topologically protected synchronization in networks
I Proof of Eq. (10) of MP
The left inequality in Eq. (10) of MP is well known. As for the right inequality consider the Laurent series for $\sin(\theta)$
$-\sin(\theta)=-\sum_{k=0}^{\infty}\frac{\theta^{2k+1}}{(2k+1)!}(-1)^{k}=-\theta+\frac{\theta^{3}}{3!}-\frac{\theta^{5}}{5!}+\frac{\theta^{7}}{7!}-\ldots.$ (S1)
Applying Eq. (S1) with $0\leq\theta<1$ we get
$-\sin(\theta)\leq-\theta+\frac{\theta}{3!}+\frac{\theta}{7!}+\ldots=-\theta(1-\gamma).$ (S2)
II Proof that $\lambda<0$ and $0<\Theta_{0}<1$ imply $0<\Theta<1$, $\forall t>0$
We limit the proof to the case of equal frequencies. Let us start from Eq. (13) of MP that here we rewrite for convenience:
${\Theta}_{0}e^{-J\mu t}<{\Theta}<{\Theta}_{0}e^{J\lambda t},$ (S3)
which is valid for $t\in[0,t_{1})$ under the hypothesis that $0<{\Theta}_{0}<1$. Recall that $0,t_{1})$ is defined as the interval where the solution satisfies $0<\Theta(t)<1$. Let us observe that the Kuramoto-like model, Eq. (1) of MP, implies that, for any $i=1,\ldots,N$, both $\theta_{i}(t)$ and $\dot{\theta}_{i}(t)$ are continuous functions of time in the interval $0,\infty)$ (they are actually class $C^{\infty}[0\,+\infty)$ but for us it suffices to observe that they are — at least — class $C^{1}[0\,+\infty)$). In turn, this implies the continuity of both the phase difference $\Theta(t)$ and its time derivative $\dot{\Theta}(t)$.
Let us suppose that, by contradiction, $t_{1}<\infty$. This implies that it is either $\Theta(t_{1})=1$ or $\Theta(t_{1})=0$. We shall see in a moment that both possibilities have to be discarded.
If we suppose the first possibility, $\Theta(t_{1})=1$, from the upper bound in Eq. (S3) it follows that, for any $0<\epsilon<t_{1}$, by using $\lambda<0$, we have
$\Theta(t_{1})-\Theta(t_{1}-\epsilon)>1-\Theta_{0}e^{J\lambda(t_{1}-\epsilon)}>1-\Theta_{0}>0.$ (S4)
The above expression would imply that $\Theta(t)$ is discontinuous at $t_{1}$.
If we suppose the second possibility, $\Theta(t_{1})=0$, then, from the lower bound in Eq. (S3) it follows that, for any $0<\epsilon<t_{1}$,
$\Theta(t_{1})-\Theta(t_{1}-\epsilon)<-\Theta_{0}e^{-J\mu(t_{1}-\epsilon)}.$ (S5)
For $t_{1}$ finite, the above expression would imply that $\Theta(t)$ is discontinuous at $t_{1}$. The proof is complete.
III Other examples of groups of topologically equivalent nodes
Depending on the relation between $k^{(\mathrm{IN})}$ and $N^{\prime}$, a group of $N^{\prime}$ topologically equivalent (TE) nodes can be realized in several ways. Fig. 1 of the main paper (MP) included cases with $N^{\prime}=2,3$ and $N^{\prime}=4$. Here, Fig. S1 shows all possible cases with for $N^{\prime}=2$ to $5$. The synchronization scenario and the topological condition (TC) run as follows:
  • When the nodes of the group have no connections among each other (Panels a, c, e, m), there is no stable attractor but, in the case of equal frequencies, the dynamics of such TE oscillators is governed by the same identical equation so that, over time, each oscillator tracks exactly the other
  • On the base of the cases that we were able to calculate explicitly so far, when the nodes of the group form a fully connected (FC) graph (like Panels b, d, l, o), the sufficient TC for synchronization of the group is given by $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$ (Panels b, d. l) or $k^{(\mathrm{OUT})}<k^{(\mathrm{IN})}$ (Panel o), depending on the case.
  • Similarly, when the group forms a disjoint union of $m$ FC graphs (Panels f, g, h), the sufficient TC for the individual synchronization of each component is $k^{(\mathrm{OUT})}\leq\min_{l\in\{1,\ldots,m\}}k_{l}^{(\mathrm{IN})}$, or $k^{(\mathrm{OUT})}\leq\min_{l\in\{1,\ldots,m\}}k_{l}^{(\mathrm{IN})}$, depending on the case, where $k_{l}^{(\mathrm{IN})}$ is the internal degree of the $l$-th FC component.
  • When the nodes of the group form instead regular polygons, unless special initial conditions are chosen, the chance for synchronization decreases or ceases. For example, for the case $N^{\prime}=4$ (Panel i) $k^{(\mathrm{OUT})}$ can be at most 1, while for the case $N^{\prime}=5$ (Panel n), $k^{(\mathrm{OUT})}$ must be 0.
Figure S1: All possible groups of $N^{\prime}$ TE nodes for $N^{\prime}=2$ to 55 (top to bottom). Here we focus only on the concept of topological equivalence within the group, while the topological equivalence with respect to the neighbors of the group — indicated by the surrounding black corona — is assumed.
IV Derivation of the TC for $N^{\prime}>2$
[…]
The resulting bounding ODE can be written vectorially:
$\dot{\bm{{\Theta}}}\leq\bm{{\Omega}}+J\bm{B}\cdot\bm{{\Theta}},$ (S6)
where $\bm{{\Theta}}=({\Theta}_{2,1},{\Theta}_{3,2},\ldots,{\Theta}_{N^{\prime},N^{\prime}-1})^{T}$, $\bm{{\Omega}}=({\Omega}_{2,1},{\Omega}_{3,2},\ldots,{\Omega}_{N^{\prime},N^{\prime}-1})^{T}$, and $\bm{B}$ is a $(N^{\prime}-1)\times(N^{\prime}-1)$ suitable matrix that depends on the parameters $N^{\prime}$, $k^{(\mathrm{OUT})}$, $k^{(\mathrm{IN})}$, $\gamma$, and on the specific way in which the TE nodes are arranged among each other. Let $\lambda_{i}$ and $\bm{u}_{i}$ be the eigenvalues and normalized eigenvectors of $\bm{B}$, respectively. The coefficients $c_{i}(t)=\bm{{\Theta}}\cdot\bm{u}_{i}$ satisfy the decoupled system $dc_{i}(t)/dt\leq\bm{{\Omega}}\cdot\bm{u}_{i}+J\lambda_{i}c_{i}(t)$. As a consequence, each component of the vector $\bm{{\Theta}}$ remains bounded if all the eigenvalues of $\bm{B}$ are negative.
Below we consider FC cases and and then regular polygons.
IV.1 $N^{\prime}=3$ FC (Panel d of Fig. S1)
Here we show the bounding solution in full detail. We have the following ODE system
$\frac{\dot{{\Theta}}_{2,1}}{J}\leq\frac{{\Omega}_{2,1}}{J}-2\sin\left({\Theta}_{2,1}\right)+\sin\left({\Theta}_{3,2}\right)-\sin\left({\Theta}_{3,1}\right)+h_{2,1}^{\prime}{\Theta}_{2,1},$
$\frac{\dot{{\Theta}}_{3,2}}{J}\leq\frac{{\Omega}_{3,2}}{J}-2\sin\left({\Theta}_{3,2}\right)+\sin\left({\Theta}_{2,1}\right)-\sin\left({\Theta}_{3,1}\right)+h_{3,2}^{\prime}{\Theta}_{3,2}.,$
to be solved with the constrain
${\Theta}_{3,1}={\Theta}_{3,2}+{\Theta}_{2,1}.$ (S7)
By using the three bounds explained previously we get the following matrix
$\bm{B}=\begin{bmatrix}k^{(\mathrm{OUT})}-3(1-\gamma)&\gamma\\ \gamma&k^{(\mathrm{OUT})}-3(1-\gamma)\end{bmatrix}$, (S8)
to be used in Eq. (S6) with the vectors $\bm{{\Theta}}=({\Theta}_{2,1},{\Theta}_{3,2})^{T}$ and $\bm{{\Omega}}=({\Omega}_{2,1},{\Omega}_{3,2})^{T}$. The eigenvalues of $\bm{B}$ are
$\lambda_{1}=k^{(\mathrm{OUT})}-3+4\gamma,$ (S9)
$\lambda_{2}=k^{(\mathrm{OUT})}-3+2\gamma,$ (S10)
with normalized eigenvectors $\bm{u}_{1}=(1,1)^{T}/\sqrt{2}$ and $\bm{u}_{2}=(1,-1)^{T}/\sqrt{2}$, respectively.
We conclude that, for a group of $N^{\prime}=3$ TE FC nodes, a sufficient TC for synchronization is $k^{(\mathrm{OUT})}-3+4\gamma<0$, which, on observing that $0<4\gamma<1$, amounts to $k^{(\mathrm{OUT})}\leq 2$, or $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$, where we have made use of the fact that, here, $k^{(\mathrm{IN})}=2$. By integrating the two normal modes $c_{1}(t)$ and $c_{2}(t)$, we obtain the following bounding solution
${\Theta}_{2,1}\leq\frac{\Omega_{2,1}+\Omega_{3,2}}{2J|\lambda_{1}|}+\frac{\Omega_{2,1}-\Omega_{3,2}}{2J|\lambda_{2}|}+\left(-\frac{\Omega_{2,1}+\Omega_{3,2}}{2J|\lambda_{1}|}+\frac{c_{1}(0)}{\sqrt{2}}\right)\exp\left(-J|\lambda_{1}|t\right)+\left(-\frac{\Omega_{2,1}-\Omega_{3,2}}{2J|\lambda_{2}|}+\frac{c_{2}(0)}{\sqrt{2}}\right)\exp\left(-J|\lambda_{2}|t\right)$ (S11)
${\Theta}_{3,2}\leq\frac{\Omega_{2,1}+\Omega_{3,2}}{2J|\lambda_{1}|}-\frac{\Omega_{2,1}-\Omega_{3,2}}{2J|\lambda_{2}|}+\left(-\frac{\Omega_{2,1}+\Omega_{3,2}}{2J|\lambda_{1}|}+\frac{c_{1}(0)}{\sqrt{2}}\right)\exp\left(-J|\lambda_{1}|t\right)-\left(-\frac{\Omega_{2,1}-\Omega_{3,2}}{2J|\lambda_{2}|}+\frac{c_{2}(0)}{\sqrt{2}}\right)\exp\left(-J|\lambda_{2}|t\right),$ (S12)
where $c_{1}(0)=\bm{{\Theta}}(t=0)\cdot\bm{u}_{1}$ and $c_{2}(0)=\bm{{\Theta}}(t=0)\cdot\bm{u}_{2}$ are the projections of the initial conditions onto the eigenvectors. Note that, for $J$ sufficiently large, all the above procedure turns out to be consistent with the required bound $0<{\Theta}_{i,j}<1/2$ for any $t$. More precisely, by using $|\lambda_{i}|>1/2$, from Eqs. (S11)-(S12) it follows that a sufficient condition for ${\Theta}_{i,j}<1/2$ is $J>4\max_{i,j}|{\Omega}_{i,j}|$. On the other hand, by using $|\lambda_{2}|>|\lambda_{1}|$ together with $\Omega_{2,1}\geq 0$ and $\Omega_{3,2}\geq 0$, it follows also that $0<{\Theta}_{i,j}$.
IV.2 $N^{\prime}=4$ FC (Panel l of Fig. S1)
Here, the ODE system must be solved with the following constrains
${\Theta}_{4,1}={\Theta}_{4,2}+{\Theta}_{3,2}+{\Theta}_{2,1},$ (S13)
${\Theta}_{4,2}={\Theta}_{4,3}+{\Theta}_{3,2},$ (S14)
${\Theta}_{3,1}={\Theta}_{3,2}+{\Theta}_{2,1}.$ (S15)
By using the three bounds we obtain the following matrix
$\bm{B}=\begin{bmatrix}k^{(\mathrm{OUT})}-4(1-\gamma)&2\gamma&\gamma\\ \gamma&k^{(\mathrm{OUT})}-5(1-\gamma)&\gamma\\ \gamma&2\gamma&k^{(\mathrm{OUT})}-4(1-\gamma)\end{bmatrix}$. (S16)
The eigenvalues of $\bm{B}$ are
$\lambda_{1}=k^{(\mathrm{OUT})}-4+3\gamma,$ (S17)
$\lambda_{2}=k^{(\mathrm{OUT})}-\frac{9}{2}+5\gamma-\frac{1}{2}\sqrt{1+16\gamma^{2}},$ (S18)
$\lambda_{3}=k^{(\mathrm{OUT})}-\frac{9}{2}+5\gamma+\frac{1}{2}\sqrt{1+16\gamma^{2}}.$ (S19)
The largest of these three eigenvalues is $\lambda_{3}$, which, by using the fact that here is $k^{(\mathrm{IN})}=3$, and the explicit value of $\gamma=0.1668651044\ldots$, gives
$\lambda_{3}=k^{(\mathrm{OUT})}-k^{(\mathrm{IN})}-\delta,\quad\delta\simeq 0.0646.$ (S20)
We conclude that for a group of $N^{\prime}=4$ FC nodes, a sufficient TC for synchronization is $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$.
IV.3 $N^{\prime}=5$ FC (Panel o of Fig. S1)
Here, the ODE system must be solved with the following constrains
${\Theta}_{5,1}={\Theta}_{5,4}+{\Theta}_{4,3}+{\Theta}_{3,2}+{\Theta}_{2,1},$ (S21)
${\Theta}_{5,2}={\Theta}_{5,4}+{\Theta}_{4,3}+{\Theta}_{3,2},$ (S22)
${\Theta}_{4,1}={\Theta}_{4,3}+{\Theta}_{3,2}+{\Theta}_{2,1},$ (S23)
${\Theta}_{4,2}={\Theta}_{4,3}+{\Theta}_{3,2},$ (S24)
${\Theta}_{3,1}={\Theta}_{3,2}+{\Theta}_{2,1}.$ (S25)
By using the three bounds we obtain the following matrix
$\bm{B}=\begin{bmatrix}k^{(\mathrm{OUT})}-5(1-\gamma)&3\gamma&2\gamma&\gamma\\ \gamma&k^{(\mathrm{OUT})}-5(1-\gamma)&2\gamma&\gamma\\ \gamma&2\gamma&k^{(\mathrm{OUT})}-5(1-\gamma)&\gamma\\ \gamma&2\gamma&3\gamma&k^{(\mathrm{OUT})}-5(1-\gamma)\end{bmatrix}$. (S26)
The eigenvalues of $\bm{B}$ are
$\lambda_{1}=k^{(\mathrm{OUT})}-5+3\gamma,$ (S27)
$\lambda_{2}=k^{(\mathrm{OUT})}-5+4\gamma,$ (S28)
$\lambda_{3}=k^{(\mathrm{OUT})}-5(1-\gamma)+\frac{3}{2}\gamma-\frac{1}{2}\sqrt{41}\gamma,$ (S29)
$\lambda_{4}=k^{(\mathrm{OUT})}-5(1-\gamma)+\frac{3}{2}\gamma+\frac{1}{2}\sqrt{41}\gamma.$ (S30)
The largest of these eigenvalues is $\lambda_{4}$, which, by using the fact that here is $k^{(\mathrm{IN})}=4$, and the explicit value of $\gamma$, gives
$\lambda_{4}=k^{(\mathrm{OUT})}-k^{(\mathrm{IN})}+\delta^{\prime},\quad\delta^{\prime}\simeq 0.51214.$ (S31)
We conclude that for a group of $N^{\prime}=5$ FC nodes, a sufficient TC for synchronization is $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}-1$.
IV.4 General FC case
Finding a bound for the largest eigenvalue of $\bm{B}$ in the general FC case remains a formidable task. In general, we expect that the maximal eigenvalue $\bm{B}$ scales as $k^{(\mathrm{OUT})}-N^{\prime}(1-\gamma)+\alpha N^{\prime}=k^{(\mathrm{OUT})}-\left(k^{(\mathrm{IN})}+1\right)\left(\gamma+\alpha\right)$, where $\alpha=\mathop{O}(1)$. This guess needs further demanding analysis.
IV.5 Regular polygon $N^{\prime}=4$ (Panel i of Fig. 1)
As shown in Fig. S1, $N^{\prime}=4$ provides the first situation where the $N^{\prime}$ TE nodes can be arranged in more than one way. Besides the FC case, where $k^{(\mathrm{IN})}=3$, there are the configurations where $k^{(\mathrm{IN})}=1$ or $k^{(\mathrm{IN})}=2$. In the former case, the subgraph of the $N^{\prime}$ TE nodes with their intra-links is a disconnected one and corresponds to the union of two disconnected pairs of TE nodes, for which we have already seen the TC (for example ${\Theta}_{1,2}=0$ and ${\Theta}_{3,4}=0$): $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$. In the latter case, the four TE nodes form a square and the obtained ODE system is no longer totally decoupled. After using the constrain $\Theta_{4,1}=\Theta_{4,3}+\Theta_{3,2}+\Theta_{2,1}$, for the bounding ODE system of the independent variables ${\Theta}_{2,1},{\Theta}_{3,2},{\Theta}_{4,3}$, we get following matrix
$\bm{B}=\begin{bmatrix}k^{(\mathrm{OUT})}-3(1-\gamma)&\gamma&\gamma-1\\ 1&k^{(\mathrm{OUT})}-2(1-\gamma)&1\\ \gamma-1&\gamma&k^{(\mathrm{OUT})}-3(1-\gamma)\end{bmatrix}$. (S32)
The eigenvalues of $\bm{B}$ are
$\lambda_{1}=k^{(\mathrm{OUT})}-2+2\gamma,$ (S33)
$\lambda_{1}=k^{(\mathrm{OUT})}-3+3\gamma-\sqrt{1+\gamma^{2}},$ (S34)
$\lambda_{1}=k^{(\mathrm{OUT})}-3+3\gamma+\sqrt{1+\gamma^{2}}.$ (S35)
The largest of these eigenvalues is $\lambda_{3}$, which, by using the fact that here is $k^{(\mathrm{IN})}=2$, and the explicit value of $\gamma$, gives
$\lambda_{3}=k^{(\mathrm{OUT})}-2+\delta=k^{(\mathrm{OUT})}-k^{(\mathrm{IN})}+\delta^{\prime\prime},\quad\delta^{\prime\prime}\simeq 0.5144.$ (S36)
We conclude that for this group, a sufficient TC for synchronization is $k^{(\mathrm{OUT})}\leq 1$.
IV.6 Regular polygon $N^{\prime}=5$ (Panel n of Fig. 1)
After using the constrain $\Theta_{5,1}=\Theta_{5,4}+\Theta_{4,3}+\Theta_{3,2}+\Theta_{2,1}$, for the bounding ODE system of the independent variables ${\Theta}_{2,1},{\Theta}_{3,2},{\Theta}_{4,3},{\Theta}_{5,4}$, we get following matrix
$\bm{B}=\begin{bmatrix}k^{(\mathrm{OUT})}-3(1-\gamma)&\gamma&\gamma-1&\gamma-1\\ 1&k^{(\mathrm{OUT})}-2(1-\gamma)&1&0\\ 0&1&k^{(\mathrm{OUT})}-2(1-\gamma)&1\\ \gamma-1&\gamma-1&\gamma&k^{(\mathrm{OUT})}-3(1-\gamma)\end{bmatrix}$. (S37)
The eigenvalues of $\bm{B}$ are
$\lambda_{1}=k^{(\mathrm{OUT})}-\frac{5}{2}-\frac{\sqrt{5}}{2}+2\gamma,$ (S38)
$\lambda_{2}=k^{(\mathrm{OUT})}-\frac{5}{2}+\frac{\sqrt{5}}{2}+2\gamma,$ (S39)
$\lambda_{3}=k^{(\mathrm{OUT})}-\frac{5}{2}+3\gamma-\frac{\sqrt{5-4\gamma+4\gamma^{2}}}{2},$ (S40)
$\lambda_{4}=k^{(\mathrm{OUT})}-\frac{5}{2}+3\gamma+\frac{\sqrt{5-4\gamma+4\gamma^{2}}}{2}.$ (S41)
The largest of these eigenvalues is $\lambda_{4}$, which, by using the explicit value of $\gamma$, gives
$\lambda_{4}=k^{(\mathrm{OUT})}-\delta^{\prime\prime\prime},\quad\delta^{\prime\prime\prime}\simeq 0.9453.$ (S42)
We conclude that for this group, a sufficient TC for synchronization is $k^{(\mathrm{OUT})}=0$.
V Simulations
The simulations reported in MP have been realized by means of the Runge-Kutta methodNumerical_Recipes_SM at fourth-order with step parameter $h=0.01$. Concerning the precision of the method, each choice of initial conditions has been accepted or not with the following criterion. If for natural frequencies all equal and $J=1$ the simulation provides global synchronization (which is an exact result) the choice is accepted, otherwise is rejected.
As for the simulations reported in Fig. 2 of MP, the initial conditions and the natural frequencies, respectively, are:
$\theta_{1}=1.2;\theta_{2}=1.5;\theta_{3}=1.6;\theta_{4}=3.2;\theta_{5}=3.8;\theta_{6}=4.1;\theta_{7}=4.2;\theta_{8}=5.3;\theta_{9}=5.4;\theta_{10}=5.5;\theta_{11}=5.6;\theta_{12}=5.7;\theta_{13}=6.5;\theta_{14}=6.6;\theta_{15}=7.4;\theta_{16}=7.5;\theta_{17}=7.6;\theta_{18}=7.9;\theta_{19}=8.0;\theta_{20}=8.2;\theta_{21}=8.3;\theta_{22}=8.4$, and $\omega_{1}=2.1,\omega_{2}=2.2,\omega_{3}=2.3,\omega_{4}=4.5,\omega_{5}=3.5,\omega_{6}=2.6,\omega_{7}=2.7,\omega_{8}=1.1,\omega_{9}=1.2,\omega_{10}=1.3,\omega_{11}=1.4,\omega_{12}=1.5,\omega_{13}=2.5,\omega_{14}=0.8,\omega_{15}=0.5,\omega_{16}=0.7,\omega_{17}=1.5,\omega_{18}=0.8,\omega_{19}=3.0,\omega_{20}=3.1,\omega_{21}=3.2,\omega_{22}=3.3$.
Here, Fig. S2 shows the same Fig. 2 of MP in further detail by means of four increasing zooms. In Fig. S3 we show also the rotating numbers for the same system of Fig. 2 of MP.
Figure S2: Four images of Fig. 2 of MP over several scales.
Figure S3: Rotating numbers for the same system of Fig. 2 of MP.
As for the simulations reported in Fig. 3 of MP, the initial conditions and the natural frequencies, respectively, are:
$\theta_{1}=1.3;\theta_{2}=1.5;\theta_{3}=2.1;\theta_{4}=3.2;\theta_{5}=3.7;\theta_{6}=4.1;\theta_{7}=5.2;\theta_{8}=5.3;\theta_{9}=5.4;\theta_{10}=5.5;\theta_{11}=5.6;\theta_{12}=6.2;\theta_{13}=6.5;\theta_{14}=6.6;\theta_{15}=7.4;\theta_{16}=7.5;\theta_{17}=7.6;\theta_{18}=7.9;\theta_{19}=8.0;\theta_{20}=8.8;\theta_{21}=9.1;\theta_{22}=9.2$, and $\omega_{1}=1.8,\omega_{2}=2.4,\omega_{3}=3.2,\omega_{4}=4.5\omega_{5}=3.5,\omega_{6}=1.6,\omega_{7}=2.9,\omega_{8}=1.1\omega_{9}=1.2,\omega_{10}=1.8,\omega_{11}=1.3,\omega_{12}=3.0,\omega_{13}=2.5,\omega_{14}=0.8,\omega_{15}=1.1,\omega_{16}=1.2,\omega_{17}=1.5,\omega_{18}=0.8,\omega_{19}=3.2,\omega_{20}=3.5,\omega_{21}=2.9,\omega_{22}=2.8$.
Finally, we have performed other tests confirming that the number of convergent lines in the rotating number graphs seems being related to the number of TE groups of the system. Specifically, by drawing 20 random samples each having the 22 frequencies $\omega_{i}$ drawn uniformly at random in the range $[0,5]$, with $J=1$, for the number of converging lines we have obtained the mean value 4.65 with the following distribution: 2,2,3,3,4,4,4,4,4,4,5,5,5,6,6,6,6,6,7,7.